Home‎ > ‎Reports‎ > ‎Crew098‎ > ‎

day08


Crew 98 Log Book for January 22, 2011

Photos of the day





Commander's Check-In Report

Virgiliu Pop Reporting

Crew Physical Status: Excellent

Time Departed/Returned from EVA: Several EVAs performed

Brief Narrative of Field Mission Results: Solar clock partially built; reconnaissance mission.

EVA Data/Interpretations: As per the EVA report

Engineering/Hab Maintenance: Another visit from DG, the friendliest Martian ever. The commander is fixing some EVA helmets and removing the old microphone system.

Report Transmission Schedule: Intended [for all reports to be sent] by 8 PM.

Plans for Tomorrow: Will be decided in the morning.

Inventory: -

Support Requested: DG is taking care of the plumbing

Miscellaneous: -


Commander's Report

Virgiliu Pop Reporting

Yesterday night, after everybody duly sent their reports, Haritina (yes, the same Haritina I mention in every report of mine - she is multifunctional, and could work as a crew-of-one) took us outside for yet another stargazing session. We saw the amazing Milky Way prior to the moonrise - this is how the sky looks on Mars, as the Red Planet has only two tiny captured asteroids, Phobos and Deimos, as moons, then the commander returned to the Hab for the comm window (ah, the joys of duty!). Then, after a spiritual movie courtesy of Dragos (who is enjoying the desert as another step in his journey of self discovery), we went to bed to the sound of tiny steps (courtesy of Jerry the Martian).

The early (as in 'RoMars Early') morning was uneventful and then the crew took turns at doing EVAs for building the solar clock. Captain Cook then assisted Dragos in writing his daily dispatch, fixed a quick soup (actually, two types: cheddar broccoli and clam chowder).

Afterwards Betty, Haritina and Dragos performed the main EVA of the day, while the captain fixed some EVA helmets. At their return, the same captain (who was busy today) cooked Martian Curry with potatoes, and wrote his report, eager to see yet again the amazing Martian sky.

Engineering Report

Haritina Mogosanu Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes, running all day

Honey (Gas Generator) (Used? Yes/No): Not used

Battery Charge Level:?
29.5 V everstart
34.6 V armor plate
40.7 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator run all day ok


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.5

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.35

Gasoline Tank (Full/0.75/0.5/0.25/Empty):1 containers full x 5 gallons 1 half

Notes/Comments:


Oil: not checked DG said to leave it to him

Kitty Oil Quantity (# of Quarts in storage):2.5 gal

Honey Oil Quantity (# of Quarts in storage):3 gal

ATV Oil Quantity (# of Quarts in storage):on the engineering work bench
2 x 946 ml

Car Oil Quantity (# of Quarts in storage):where is it?


Water Status:

Outside Potable Water Tank Level (inches from bottom): 11

Trailer Potable Water Tank Level (inches from bottom): almost full

Hab Potable Water Tank Level (inches from bottom): 8

Potable Water Meter Reading: 411762

Notes/Comments:


GreenHab:

Condition of plants in Tank 1: plants dead please advise

Condition of plants in Tank 2: very poor

Condition of plants in Tank 3: improving

Greenhab notes/ comments:
Please send a new shipment if plants for aquatic tank 1 as they were all dead. Also, aquatic tank 2 needs some reinforcements.


Observatory:

Telescope Used? (Yes/no) yes - please see additional report below

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no):No
Oil Checked (yes/no):No
Tire Status:not checked
Odometer at end of day:-123188
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: y/n y
Oil Checked: y/n y
Fuel Consumed: filled
Tire Status:checked
Comments:

ATV 2 (Spirit):
Used: y/n y
Oil Checked: y/n y
Fuel Consumed:replaced
Tire Status:checked
Comments:

ATV 3 (Viking I):
Used: y/n n
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status: not checked
Comments: started

ATV 4 (Viking II):
Used: y/n y
Oil Checked: y/n y
Fuel Consumed: replaced
Tire Status: checked
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 70

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 60


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: all ok


General Engineering
1. Notes/Comments:
DG was here again and made major repairs to the plumbing and the insulation as you may know.


Supplementary - Musk Observatory Status Report

1. Dome
Engine not functional, the box in which the engine is located needs reinforcing or replacement;

2. Telescope
- Fully functional, I practiced polar alignment last night all night and I was happy with the result.
- Fit to use for visual astronomy albeit there was a small error that could be easily corrected in the finderscope provided the person who uses it knows the sky well.
- Both slew and slow motions work perfectly.

2. Robofocus is working very well

3. Alignment and safety:
3.1 Pier - offset alignment with the telescope mount (A, B, C)
3.2 Pier-to-telescope Mount - has one screw broken (image D) and DG who came today was not able to remove it in order to replace it. This makes the telescope wobbly on the altitude axis. This is not acceptable for the safety of the telescope not to mention imaging with it. DG reinforced the other two screws. We took the telescope down and tried to look inside the pier but we decided that it was not safe to open the cylinder with the controls without knowing what was in there. We put the telescope back. (Note: The telescope is designed to be transportable - and it could sit very well on a tripod too - therefore I considered the risk of removing it from the pier acceptable as I have handled these telescopes before.)

There are two options to fix this without opening it, either to cut the screw or light weld a reinforcement bar and DG he said that he could do that. Please advise ASAP, it would be great to take advantage of me being here and get the telescope issue out of the way, leaving it fully functional - to include astrophotography. Besides it would make me happy too as I would have the chance to be useful :)

3.3 Leveling - telescope is now leveled (Image "F" before, Image "G" after). To do that we screwed the bolts at the bottom of the pier.

PDF of telescope mount images

EVA 6 Report

Haritina Mogosanu Reporting

Time: 11.00-11.20

Crew Members: Haritina Mogosanu, Dragos Bratasanu

Site Location: N 38.405278 (this is the site for the sundial) W 110.78333

Check where the Sun's shadow falls at 1100 hours and align it with the stone already placed there since yesterday Transit Mode: on foot

Objectives:

1. checked that 1100 hours aligned with the stone for accurate time

Results: Stone aligned

Lessons Learned: it takes at least half hour to prepare and dress up for an EVA

EVA 7 Report

Haritina Mogosanu Reporting

Time: 1200-1220

Crew Members: Haritina Mogosanu, Beatrice Gilea

Site Location: N 38.405278 (this is the site for the sundial) W 110.78333

Checked if the shadow cast by the Sun at meridian (1200 hour) aligned with the stone that points North

Transit Mode: on foot

Objectives: 1. checked that the shadow cast at 1200 hours aligned with the stone that points North

Results: Stone aligned perfectly
Another photo of stone alignment

Lessons Learned: It takes a lot of dedication to have an EVA every hour


EVA 8 Report

Haritina Mogosanu Reporting

Time: 1300-1320

Crew Members: Iulia Jivanescu, Beatrice Gilea

Site Location: N 38.405278 (this is the site for the sundial) W 110.78333

Checked that the shadow cast by the Sun at 1300 hours aligned with the stone already placed there

Transit Mode: on foot

Objectives:

1. checked that the shadow cast at 1300 hours aligned with the stone already placed in the dial

Results: Stone realigned by hand

Lessons Learned: Is very hard to move big stones on a windy Martian afternoon

EVA 9 Report

Haritina Mogosanu Reporting

Time: 1400-1420

Crew Members: Iulia Jivanescu, Beatrice Gilea

Site Location: N 38.405278 (this is the site for the sundial) W 110.78333

Checked time 1400 hours, aligned stone already placed there

Transit Mode: on foot

Objectives:

1. checked that the shadow cast at 1400 hours aligned with the stone already placed in the dial

Results: Stone realigned by hand

Lessons Learned: Is very hard to move big stones on a windy Martian afternoon

EVA 10 Report

Haritina Mogosanu Reporting

Time: 1630-1800

Crew Members: Haritina Mogosanu, Dragos Bratasanu, Beatrice Gilea

Site Location: Origin N 38.405278 W 110.78333

Reconaissance mission - map here ... and here

Transit Mode: ATV

Objectives:

1. Reconnaissance for the photo shoot session tomorrow.

Results:Found some more breathtaking scenery

Lessons Learned: Very hard to drive with the backpack on.

EVA 11 Report

Haritina Mogosanu Reporting

Time: 1930-2030

Crew Members: Virgil Pop, Haritina Mogosanu, Dragos Bratasanu, Iulia Jivanescu, Beatrice Gilea

Site Location: 38.406496, 110.791943

First night EVA mission to stargazing

Transit Mode: on foot

Objectives:

1. learning how to navigate by the stars

Results: Covered: Constellations and asterisms, constellation identification. RA, Dec, first meridian of Right Ascension, Milky Way, ecliptic, celestial equator, landscape astrophotography basics,

Lessons Learned: only astronomical :)

Journalist Report

Dragos Bratasanu Reporting

It seems that most of the crew members felt like this is a weekend even on Mars. They tried to sleep in but unfortunately the mice in the attic had other plans, waking us up at 8. We still haven't figured out how many of them are up there but it sounded like they were playing football with some food.

Haritina took each of us in short EVAs near the hab to build the solar dial. This will be done tomorrow. This will resemble the ancient dacian (people that lived on the present territory of Romania over 2000 years ago).

In the evening we went on a EVA to discover new places for scientific purposes and photo shootings.

Our commander is currently pursuing his cooking chores and the rest of the team are writing reports of the day

Cook's Report

Virgiliu Pop Reporting

1a) Today was
- a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?
- No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.
- Morning meal: tortillas (the penultimate pack) with peanut butter and jam – the crew did not want scrambled eggs. Beverages: coffee and tea.
- Lunch: Clam chowder (pack + cans of clams) for three crew members; broccoli and cheddar soup for the other two. It is true what it is said about seafood: you either love it or hate it.
- Evening meal: Alien Life Form Curry with spicy potatoes

3) List any main meal foods not finished by the crew
- None

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.
- Alien Life Form Curry with Spicy Fried Potatoes
Take freeze dried roast beef dice, dehydrated onions, dehydrated broccoli, dehydrated mushrooms and dehydrated potatoes and hydrate them in separate mugs.

Heat 1-2 tablespoons of ghee, and fry the onions in it; add curry powder and soy sauce to taste, and further fry the mix. Add the broccoli and mushrooms and let it cook for a few minutes to blend the flavors.

Fry the rehydrated potatoes with oregano and red pepper until crispy, pretending they are french fries.

You pray that the crew will survive, then you are surprised they are very very happy with the meal. If you are called Virgil, you add loads and loads of Tabasco sauce, Worcestershire sauce, and sesame oil. In fact, if you are called Virgil, you would add Tabasco to everything.

5) Comments/questions:
- Why did the chicken cross the road? Andwer: to get freeze-dried and shipped to Mars.